Shift summary for 19 Sept. 1999 evening.

Nickolas Solomey (solomey@hep.uchicago.edu)
Sun, 19 Sep 1999 23:57:37 -0500

Shift Summary for eve shift, 19/Sept/99

On Shift: Ricardo, Nick

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: 7.25
Hours of E799 data (special runs not included): 6.75? (good data yet??)
Hours of special runs (laser scans, muon runs...): (6.75 or 0) depends on
answer to previous question.

Accounting of lost time: Accelertor off 1 hr and run aborted with tape
problems.

========================================================================
Runs ended this shift:
14633 aborted
14634 failed to start due to hang on ktev3 tape, concluded tape has no eot
14635 still running.

Tapes brought to FCC this shift: none, the old set of tapes left here
for experts because they might not have an eof or eot mark, someone
should decide what to do with them.

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: mostly data taking, but uncertain if this is good
data taking yet.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:30 Inherit shift, no beam due to Booster problems
14633 15:52 Start run, a new level 2 for trigger 16 by Arun.
27 16:29 Beam returns
16:35 Conclusion is that the level 2 for trigger 16
is busted, even after removing the l2 requirments
nothing from the counting l1 is getting threw
the l2, so it is a hardware problem.
39 Accelerator miss one spill with no beam.
51 Run aborted
14634 17:15 Could not get tape drive on ktev3 to start,
suspect no eof on tapes so we are changing to new
tapes.
14635 17:30 New tapes, starting new run after drt_fresh_start.
17:36 Data taking starts after recovery from DA
problems.
15? KTeVana errors, several kfrends needed to make it
take data again.
30-31 Have no data in these two spill because accelertor
missed those spills.
173 21:26 Accelertor lost one spill.
288 24:00 End of shift, beam still coming.

=========================================================================
Follow-up on next shift:
1)Clean out the KTeV refrigerator, before the whole colaboration gets
sick.
2)Do a muon run and flip analysis magnet afterwards.
3)Look into debugging trigger 16 which is still not working.
4)Order tonner cartridges for the laser printer, monday day shift.
5)Valerie said to answer yes at next run to update constants.
6)Jason has modifed the hcc namelist that will be loaded next run, if it
does not work page him.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey